The water line is given by the equation y=-2/3x-12 Imagine you want to put a pink flamingo lawn ornament in your backyard. You want to avoid placing it directly over the water line in case you need to excavate the line for repairs in the future. Could you place it at the point (-4,-10)?
Answer by dabanfield(803)

The water line is given by the equation y=-2/3x-12 Imagine you want to put a pink flamingo lawn ornament in your backyard. You want to avoid placing it directly over the water line in case you need to excavate the line for repairs in the future. Could you place it at the point (-4,-10)?
We need to see if (-4,-10) is a point on the line. If it is then when x = -4, y = -10. Substituting in the equation for the line we have:
y = -2/3x - 12
y = (-2/3)*-4 - 12
y = 8/3 - 12
y = 8/3 - 36/3
y = -28/3
y = -9 1/3
y is not equal to -10 so the point is not on the line.